OpenStudy (anonymous):

The cost of a ticket to an amusement park was increased by 4 dollars. The new cost is twenty-eight dollars. a. Write a one-step equation that could be solved to find the old price. Let p represent the old price. b. Solve the equation you wrote in Part a. Show all your work. c. Explain how you know that the solution you found in Part b is the correct answer.

OpenStudy (anonymous):

a. p + 4 = 28 b. p + 4 - 4 = 28 - 4 p = 24 c. Substitute 24 for p: 24 + 4 = 28 28 = 28 True

OpenStudy (anonymous):

thanks